OpenStudy (anonymous):

john has a jar filled with 150 coins made up of quarters and nickels the total value of the coins is 26.70 how many quarters and how many nickels are in the jar

OpenStudy (anonymous):

{25 q + 5 n = 2670, q + n = 150} { n = 54, q = 96 }

OpenStudy (anonymous):

John has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $26.70. Let x represent the number of quarters in the jar. Which equation can be used to find the number of quarters? A. 0.05x + 0.25(150 – x) = 26.70 B. 0.25x + 0.05(150 – x) = 26.70 C. 0.25x + 0.05(x – 150) = 26.70 D. 0.05x + 0.25(x – 150) = 26.70